OpenStudy (anonymous):

Write an equation for tangent given the period, phase shift, and vertical shift. (check answer)

OpenStudy (anonymous):

period =1/3pi , phase shift = –1/4pi, vertical shift = –5

OpenStudy (anonymous):

I know it's either (3x+-3/4pi)-5. but I'm not sure whether the 3/4 is negative or positive.